OpenStudy (anonymous):

1. Find the area of a parallelogram with base b and height h. b = 82 cm h = 16.6 cm NOTE: Area of a Parallelogram = base x height (1 point) • 1,361.2 cm2 • 137.78 cm2 • 98.6 cm2

OpenStudy (anonymous):

is it base times height?

OpenStudy (anonymous):

Your answer is A. It's quit simple they give you the formula all you have to do is fill it in. Don't get confused by the question.

OpenStudy (anonymous):

Seems very straightforward. What part do you not understand?

OpenStudy (anonymous):

1361.2 cm^2

OpenStudy (anonymous):

ur answer is A

OpenStudy (anonymous):

nikias faster than me XD

OpenStudy (anonymous):

. Find the area of the triangle. (1 point) NOTE: Area of a Triangle = 1 /2 x (base x height) • 86.1 yd2 • 344.4 yd2 • 172.2 yd2 

OpenStudy (anonymous):

P=B*H They give you base which is 82 then they give you height which is 16.6. And again you formula is p=b*h now fill them in you don't have P but you have B and H. @luvalpaca lmao thank you.

OpenStudy (anonymous):

XD yuppers

OpenStudy (anonymous):

@OpenOwlDegree the answer to your question your graph, second question is 195.3 i see is not an option on there.